Operation Summary for November 6, 2009 - November 9, 2009 |
The Integrated Luminosity for the period from 11/2/09 to 11/9/09 was 44.22 pb-1. Friday Nov. 6 The day shift began with Operations monitoring the Tevatron (TeV) store 7325 (44.63E30) and the Recycler stash (216.8E10), with the Antiproton Source stacking (18.0E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. During the day shift, at 11 AM Operations disabled the Linac beam to swap out a 400 MeV area chopper tube and repair a LCW leak on a Linac RF station (LRF3) driver tube. Operations had the Linac beam re-established by 1 PM. During the evening shift, Operations terminated store 7325 at 8 PM and put the TeV into HEP shot setup. LRF2 became problematic and operators called the Linac on-call person for assistance. A control card for a MI quadrupole power supply (M:Q107) failed; operators replaced the card. Operations established store 7326 at 9:30 PM with an average initial luminosity (AIL) of 245.8E30. Saturday Nov. 7 The midnight shift began with Operations monitoring store 7326 (152.33E30) and the stash (111.2E10), with Pbar stacking (24.6E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, LRF2 tripped off; it reset. During the day shift, Operations terminated store 7326 at 11 AM. Operations put the TeV into shot setup, but the TeV quenched at sector A4 and B0 low betas. The cause of the quench was due to a higher number of protons in the TeV. During the evening shift, Operations put the TeV into shot setup at 6 PM and established store 7331 at 7:30 PM with an AIL of 136.6E30. Sunday Nov. 8 The midnight shift began with Operations monitoring store 7331 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, BRF19 tripped off and would not reset. Operations turned on BRF1 in its place. Operators tuned the Booster and returned the NuMI beam intensity to its nominal. During the day shift, Operations terminated store 7331 at 10:30 AM. Operations put the TeV into shot setup and established store 7332 at noon with an AIL of 262.5E30. Operators tuned the Proton Source for increased efficiencies. During the evening shift, Operations investigated an intermittent Pbar controls/timing problem. Monday Nov. 9 The midnight shift began with Operations monitoring store 7332 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations terminated store 7332 at 4:01 AM and put the TeV into shot setup. Operations established store 7333 at 5:40 AM with an AIL of 248.2E30. The Plans for Monday The plans for today are to stack, stash, and store. The Run Coordinator will allow Booster and MI personnel to conduct studies. |
|
More Information Comments and Suggestions |
| last modified 11/09/2009 email Fermilab |
| Security, Privacy, Legal | |